On January 1, 2018, Brown Co. borrowed cash from First Bank by issuing a $52,000 face value, four-year term note that had an 6 percent annual interest rate. The note is to be repaid by making annual cash payments of $15,007 that include both interest and principal on December 31 of each year. Brown used the proceeds from the loan to purchase land that generated rental revenues of $24,960 cash per year. Required a. Prepare an amortization schedule for the four-year period. (Round your answers to the nearest whole dollar amount.) BROWN CO Amortization Schedule Principal Balance Cash Payments Applied toApplied to Principal B Year 2018 2019 2020 2021 on January 1 December 31 Interest Principal End of Period

b. Organize the information in accounts under an accounting equation. (Round your answers to the nearest whole dollar amount. Enter any decreases to account balances with a minus sign. Select NA if there is no effect on the Accounts Titles / Retained Earnings.) BROWN CO Effect of Events on the Accounting Equation 2018, 2019, 2020 and 2021 Liabilities Stockholders Assets Accounts Titles / + Retained Retained Earnings Event Cash Land Notes Payable Earnings 2018 12/31 12/31 Bal 2019 Beg. bal. 12/31 12/31 End. bal. 2020 Beg. bal. 12/31 12/31 End. bal. 2021 Beg. bal. 12/31 12/31 End. bal

c. Prepare an income statement, a balance sheet, and a statement of cash flows for each of the four years. Complete this question by entering your answers in the tabs below. Req C Inc Req Bal R C CF Stmt Sheet Stmt Prepare the income statement for each of the four years. (Round your answers to the nearest whole dollar amount.) BROWN cO Income Statement For the Year Ended December 31 2018 2019 2020 2021

Req C Bal Req C Inc Stmt Sheet Req C CF Stmt Prepare the balance sheet for each of the four years. (Round your answers to the nearest whole dollar amount.) BROWN Co Balance Sheet As of December 31 2018 2019 2020 2021 Assets Total assets Liabilities Total liabilities Stockholders Equity Total stockholders equity Total liabilities and stockholders equity

Req CIncReq C CF Stmti Stmt Sheet Prepare the statement of cash flows for each of the four years. (Round your answers to the nearest whole dollar amount. Amounts to be deducted should be indicated with a minus sign.) BROWN co Statement of Cash Flows For the Year Ended December 31 2018 2019 2020 2021 Cash flows from operating activities Net cash flow from operating activities Cash flows from investing activities Net cash flows from investing activities Cash flows from financing activities Net cash flows from financing activities Net change in cash Ending cash balance
Answer #1

$24,960 $52,000 Rent revenue per year Loan Amount Annual Interest rate Term Annual Payment 4 Years $15,007 Amortization table of the Loan can be created as below: Interest ExpPrinciple Paid Principle Balance on Payments (Initial Carrying Value Annual (Monthly Payment- Principle Balance End Cash December 31 Interest rate Interest Expense) of the period $40,113 $27,513 $14,157 Year $52,000 S40,113 $27,513 $14,157 $15,007 $15,007 $15,007 $15,007 $11,887 $12,600 $13,356 $14,158 12 2018 2019 2020 2021 $2,407 $1,651 15 Effect of events on the accounting equation 2018, 2019, 2020 and 2021 Event Stockholders Equity Accounting Titles Retained Earnings Retained Earnin Asset Liabilities Cash Notes Payable 2018 $52,000+ $11,887)+ $40,113+ $40,113 12,600)+ 27,513+ $27,513 $52,000 $52,000) ($15,007 52,000 31-Dec 31-Dec $3,120 $24,960 52,000 $24,960 $24,960 $24,960 2019 29 $15,007 $52,000 ($2,407 5,527 $5,527 30 31-Dec $15,007 30,014) $30.014 52,000 $49,920 S49,920 24,960 2020 34 $52,000 ($15,007) $1,651 31-Dec 31-Dec $45,021 $52,000- 14,157+ $7,178 $74,880 $74,880 $24,960 2021 $52,000 14,157 ($14,158)+ ec 15,00 31-Dec $60,028 $52,000 - $99,840Formula sheet

A B C D E F G H I J K
2
3 Rent revenue per year 24960
4 Loan Amount 52000
5 Annual Interest rate 0.06
6 Term 4 Years
7 Annual Payment 15007
8
9 Amortization table of the Loan can be created as below:
10
11 Year Principle Balance on Jan 1 Cash Payments December 31 Interest Expense (Initial Carrying Value*Annual Interest rate) Principle Paid (Monthly Payment - Interest Expense) Principle Balance End of the period
12 2018 =D4 =$D$7 =D12*$D$5 =E12-F12 =D12-G12
13 2019 =H12 =$D$7 =D13*$D$5 =E13-F13 =D13-G13
14 2020 =H13 =$D$7 =D14*$D$5 =E14-F14 =D14-G14
15 2021 =H14 =$D$7 =D15*$D$5 =E15-F15 =D15-G15
16
17
18 Effect of events on the accounting equation
19 2018, 2019, 2020 and 2021
20 Event Asset = Liabilities + Stockholders' Equity Accounting Titles /
21 Cash Land = Notes Payable + Retained Earnings Retained Earnings
22 2018
23 43101 =D4 = =D23 +
24 43101 =-D23 =-D24 = +
25 43465 =-E12 = =-G12 + =-F12
26 43465 = + =D3
27 Bal =SUM(D23:D26) =SUM(E23:E26) = =SUM(G23:G26) + =SUM(I23:I26) =SUM(J23:J26)
28 2019
29 Beg. Bal =D27 =E27 =G27 =I27 =J27
30 43465 =-E13 = =-G13 + =-F13
31 43465 = + =$D$3
32 Bal =SUM(D29:D31) =SUM(E29:E31) = =SUM(G29:G31) + =SUM(I29:I31) =SUM(J29:J31)
33 2020
34 Beg. Bal =D32 =E32 =G32 =I32 =J32
35 43465 =-E14 = =-G14 + =-F14
36 43465 = + =$D$3
37 Bal =SUM(D34:D36) =SUM(E34:E36) = =SUM(G34:G36) + =SUM(I34:I36) =SUM(J34:J36)
38 2021
39 Beg. Bal =D37 =E37 =G37 =I37 =J37
40 43465 =-E15 = =-G15 + =-F15
41 43465 = + =$D$3
42 Bal =SUM(D39:D41) =SUM(E39:E41) = =SUM(G39:G41) + =SUM(I39:I41) =SUM(J39:J41)
43